OpenStudy (gabylovesyou):

Identify the correct slope and y-intercept of the equation 5x - 3y = -6

OpenStudy (baldymcgee6):

Solve for y

OpenStudy (baldymcgee6):

and then when you want to find the y intercept let x=0

OpenStudy (baldymcgee6):

as soon as you solve for y, the slope will become apparent.

OpenStudy (gabylovesyou):

slope = -5; y-intercept = (0, -6)

OpenStudy (gabylovesyou):

slope = 5/3; y-intercept = (0, 2)

OpenStudy (baldymcgee6):

correct, when you solve for y, you will get something in the form of y= mx+b, where m is your slope and b is your y intercept!

OpenStudy (baldymcgee6):

so in this case y = 5x/3 +2 slope of 5/3 y int. (0,2)